OpenStudy (anonymous):

ans. the following: x^2-2x=3 x^2-12x=64 4x^2+4x-24=0 9x^2-12x-5=0 16x^2-8x=15

OpenStudy (anonymous):

wha we hv to find??

OpenStudy (theviper):

Every equation can be solved by the formula\[\Huge{\color{red}{-b \pm \sqrt{b^2-4ac}\over2a}}\]

OpenStudy (theviper):

Equations 1 is as follows:- \[1.)x^2-2x-3=0\]

OpenStudy (anonymous):

here b=-2, a=1, c=-3

OpenStudy (anonymous):

what is the ans.?

OpenStudy (anonymous):

for the first 1...x = 7 or -9

OpenStudy (anonymous):

@florrdeliza dont ask for answer directly...solve it urself...it will help u a lot..

OpenStudy (anonymous):

x^2-12x-64=0 x^2-16x+4x-64=0 x(x-16)+4(x-16)=0 (x+4)(x-16)=0 x=-4 or x=16

OpenStudy (anonymous):

4x^2+4x-24=0 4x^2+12x-8x-24=0 4x(x+3)-8(x+3)=0 (4x-8)(x+3)=0 x=2 or x=-3

OpenStudy (anonymous):

excusme i am solving your problems..........atleast see them.........